OpenStudy (anonymous):

Find if divergent/convergent: a(sub n)=(n^2/sqr(n^3+4n)), if convergent, find the limit.

OpenStudy (zarkon):

where are you stuck??

OpenStudy (anonymous):

how to start it <<<

OpenStudy (zarkon):

what tests do you know?

OpenStudy (anonymous):

squeeze theorm, using f(x) and take the limit

OpenStudy (anonymous):

this is what i tired to do.... f(x)=(x^2(/sqr(x^3+4x)), and take the limit, but dont know how

OpenStudy (zarkon):

\[\sqrt{n^3+4n}=\sqrt{n^3\left(1+\frac{4}{n^2}\right)}=n^{3/2}\sqrt{1+\frac{4}{n^2}}\]

OpenStudy (anonymous):

yes, i divided top/bottom by n^3, and got (1/n)/(sqr(1+(4/(n^2))))

OpenStudy (anonymous):

so it divergent because top is zero when take the limit?

OpenStudy (zarkon):

try that again...you didn't do it correctly...use what I wrote

OpenStudy (anonymous):

sqr(n)/((sqr(1+4/(n^2))) ??

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

so limit as x>inifnity, is infinty/1 ?,, that doesnt seem right

OpenStudy (zarkon):

so the top is going to infinity...the bottom to 1...that means that the limit is infinity (or does not exist..depending on what the teacher wants)

OpenStudy (anonymous):

i c... thanks, the book says Divergent, so it is infinity/not exist.

OpenStudy (zarkon):

same thing

OpenStudy (anonymous):

kk, ty for the help